# Enhancing Dynamic CT Image Reconstruction with Neural Fields and Optical Flow

### v2

@@ 4.1 Two-square phantom
For this phantom, we set the distance source-origin to 3, the distance source-detector to 5, and the detector size to 3.5 to specify the projection geometry. We consider $N_{T}=100$ frames, leading to a measurement array of dimension $64\times 100$. Measurements are further corrupted with Gaussian noise with standard deviation $\sigma=0.01$. See figure 3(b). During reconstruction, each frame is generated by evaluating the neural field at a spatial grid of resolution $64\times 64$.
@@ 4.2 Cardiac phantom
We set the same projection geometry as for the two-square phantom. For this phantom we consider $N_{T}=300$ frames, leading to a measurement array of dimension $64\times 300$. Measurements are further corrupted with Gaussian noise with standard deviation $\sigma=0.01$. See figure 4(b). During reconstruction, each frame is generated by evaluating the neural field at a spatial grid of resolution $64\times 64$.
@@ 4.3 STEMPO
In [54], it is introduced the Spatio-TEmporal Motor-Powered (STEMPO) dataset. It provides X-ray tomography data for a moving object. The dynamics of the object are controlled by a motor allowing for different sampling schemes to be performed. The phantom consists of two static objects and a square moving upwards from the bottom to the top at a constant speed.
@@ 5 Numerical Experiments
As mentioned before, in our numerical experiments, we use Tomosipo to compute the X-ray transform and its transpose. The architecture for the neural networks $u_{\theta}$ and $\mathbb{v}_{\phi}$ consists of an input layer of size 3, then Fourier feature mappings of the form $(x,t)\in\mathbb{R}^{3}\to(\Gamma_{1}(x),\Gamma_{2}(t))\in\mathbb{R}^{128}$, then three hidden layers of size 128 and we finish with an output layer of size 1 for $u_{\theta}$ and size $2$ for $\mathbb{v}_{\phi}$. We do not apply an activation function in the last layer. In particular, we do not impose the network $u_{\theta}$ to output non-negative values. The mappings are defined as $\Gamma_{1}(x):=(\sin(2\pi\mathbf{B_{x}}x),\cos(2\pi\mathbf{B_{x}}x))$ and $\Gamma_{2}(t):=(\sin(2\pi\mathbf{B_{t}}t),\cos(2\pi\mathbf{B_{t}}t))$, with the matrices $\mathbf{B_{x}}\in\mathbb{R}^{32\times 2}$ and $\mathbf{B_{t}}\in\mathbb{R}^{32\times 1}$ having non-trainable entries $(\mathbf{B_{x}})_{ij}\sim\mathcal{N}(0,\sigma_{x}^{2})$ and $(\mathbf{B_{t}})_{ij}\sim\mathcal{N}(0,\sigma_{t}^{2})$. The sinusoidal functions act element-wise. We found $\sigma_{x}=\sigma_{t}=0.1$ to give the best results for the two-square and the STEMPO datasets, and $\sigma_{x}=0.1,\sigma_{t}=0.5$ for the cardiac dataset. We try two batch settings, the full-batch $N_{b}=N_{T}$ and the mini-batch $N_{b}=1$. In the full-batch setting we optimize for 150,000 iterations, while in the mini-batch setting, we optimize for 10,000 epochs (leading to $10,000\times N_{T}$ iterations). We set the sampling rate (8) at $SR=0.1$, leading to 40,960, 122,880, and 44,100 collocation points being randomly sampled in each iteration for the two-square, cardiac, and STEMPO datasets, respectively. We use the Adam optimizer with a learning rate of $10^{-3}$ in all cases. For the grid-based approach, each subproblem, (10) and (11), are run for 2,000 iterations and this alternation is repeated 5 times. The PSNR values reported in this section are computed using the resolution during reconstruction. Since the ground truth is defined at a higher resolution in space, we downsample it using pooling average. Experiments are performed on a Nvidia Tesla V100 GPU with 16GB [55].
5.1 Unregularized versus regularized neural fields
@@
Figure 6: Ablation study on $\gamma$ for neural fields. Evolution of PSNR every 100 iterations during optimization with batch size $N_{b}=N_{T}$.
@@
Figure 7: Evolution of PSNR every 100 iterations for the unregularized ($\gamma=0$) and the proposed regularized solution $(\gamma=10^{-2})$ with neural fields.
@@
Figure 8: Varying batch size $N_{b}\in\{1,100\}$ for $\gamma=0$ and $\gamma=10^{-2}$. Evolution of PSNR against time in hours during optimization of the neural field.
@@
Figure 9: Reconstruction (left) with neural fields and error (right) of the two-square phantom for the random sampling strategy at frames 35, 55, and 75 (out of 100).
@@
In this section, we study the effect of optimizing the neural field with an explicit motion regularizer and compare it against the purely implicitly regularized solution. For this purpose, we set the regularization parameters $\alpha=\beta=0$ and perform an ablation study on $\gamma\in\{10^{-4},10^{-3},10^{-2},10^{-1},1\}$ in the three datasets with random sampling and measure its performance in terms of PSNR.
@@
We also compare the full-batch setting $N_{b}=N_{T}$ against the commonly used mini-batch setting with $N_{b}=1$, and for $\gamma\in\{0,10^{-2}\}$. We recall that for $N_{b}=1$ optimization is done for 10,000 epochs and results are displayed every 100 epochs. Results for the evolution of PSNR versus time are shown in figure 8. A batch size of 1 speeds up the optimization achieving a higher PSNR in less time. It suffers however from an unstable behaviour with the PSNR varying for almost 2 points every 100 epochs. This means we can stop at a poor reconstruction without a reliable stopping criteria.
We show the obtained reconstructions at their best PSNR and the corresponding error against the ground truth at different frames for the two-square, cardiac, and STEMPO datasets in figures 9, 11, and 11, respectively. The top rows show the full-batch regularized case $N_{b}=N_{T},\gamma=10^{-2}$; the middle rows show the mini-batch regularized case $N_{b}=1,\gamma=10^{-2}$; the bottom rows show the full-batch unregularized case $N_{b}=N_{T},\gamma=0$. A relevant advantage in the mini-batch setting is that it is more likely to capture edges in less time since more iterations can be taken, while its full-batch counterpart $N_{b}=N_{T}$ shows blurry edges. We also summarize the PSNR values obtained for the different models in table 1 highlighting the role of $\gamma$.
@@
Figure 10: Reconstruction (left) with neural fields and error (right) of the cardiac phantom for the random sampling strategy at frames 1, 122, and 250 (out of 300).
Figure 11: Reconstruction (left) with neural fields and error (right) of the STEMPO phantom for the sequential 32-degree sampling strategy at frames 1, 40, and 60 (out of 90).
5.2 Regularized neural fields versus grid-based method.
Figure 12: Neural field versus grid-based for the two-square phantom.
Figure 13: Neural field versus grid-based for the cardiac phantom.
@@
Figure 14: Comparing grid-based method against neural fields for the two-square phantom with sequential sampling. Reconstruction (left) and error (right).
Figure 15: Comparing grid-based method against neural fields for the STEMPO phantom with sequential 4-degree sampling. Reconstruction (left) and error (right).
We now compare explicitly regularized neural fields against the grid-based method outlined in section 3.2. For this approach, it was found that the choice $(\alpha,\beta,\gamma)=(10^{-3},10^{-4},10^{-3})$ led to the best results in terms of PSNR, achieving a value of 27.09. We solve the same variational problem with neural fields by choosing the same regularization parameters, in which case the PSNR achieved is 32.92. Figure 12(a) shows reconstructions at different frames. To appreciate the time evolution we also show a $x-t$ slice view in figure 12(b) at the horizontal center $y=0$, where it becomes clear that the grid-based solution struggles at capturing regularity in time.
@@
We finish our study by comparing both methods at the more challenging problem of sequential sampling. We try the sequential 9-degree sampling for the two-square phantom and the 4-degree sampling for STEMPO. For the two-square phantom, we use the regularization parameters that gave the best results in the previous section, namely, $(\alpha,\beta,\gamma)=(10^{-3},10^{-4},10^{-3})$ for the grid-based method achieving a PSNR of 22.65. For the neural field, we use $(\alpha,\beta,\gamma)=(0,0,10^{-2})$ and the method achieves a PSNR of 26.42. Results are shown in figure 15. There it can be seen that both methods perform worse than in the random sampling, but still, the neural field reconstruction is better than the grid-based one which shows large errors, for instance, in frame 60. For STEMPO we found $(\alpha,\beta,\gamma)=(10^{-4},10^{-4},10^{-1})$ to give the best reconstruction for the grid-based method with a PSNR of 14.24; for the neural field, we set $(\alpha,\beta,\gamma)=(0,0,10^{-2})$, leading to a PSNR of 15.18. Results can be seen in figure 15. In this case, both methods can capture the static part of the image but fail at the reconstruction of the moving square due to the sampling scheme.
@@ 6 Conclusion
This work considers neural fields for dynamic CT image reconstruction from finely resolved in time but severely undersampled in space measurements as commonly encountered in applications, e.g. cardiac CT. Neural fields are particularly suitable for this task: their continuity allows for a coherent representation in both time and space; their resolution independence enables us to query it at multiple frames, and, together with their differentiability, facilitate the numerical evaluation of the PDE-based motion regularizer. We studied how to enhance the neural field reconstruction by making use of the optical flow equation: constraining the neural field to this physically feasible motion meant a significant improvement with respect to the state-of-the-art both in terms of neural fields as well as traditional grid-based representations.

### v3

@@ 4.1 Two-square phantom
For this phantom, we set the distance source-origin to 3, the distance source-detector to 5, and the detector size to 3.5 to specify the projection geometry. We consider $N_{T}=100$ frames, leading to a measurement array of dimension $64\times 100$. Measurements are further corrupted with Gaussian noise with standard deviation $0.01$. See figure 3(b). During reconstruction, each frame is generated by evaluating the neural field at a spatial grid of resolution $64\times 64$.
@@ 4.2 Cardiac phantom
We set the same projection geometry as for the two-square phantom. For this phantom we consider $N_{T}=300$ frames, leading to a measurement array of dimension $64\times 300$. Measurements are further corrupted with Gaussian noise with standard deviation $0.01$. See figure 4(b). During reconstruction, each frame is generated by evaluating the neural field at a spatial grid of resolution $64\times 64$.
@@ 4.3 STEMPO
In [58], the Spatio-TEmporal Motor-Powered (STEMPO) dataset is introduced. It provides X-ray tomography data for a moving object. The dynamics of the object are controlled by a motor, allowing for different sampling schemes to be performed. The phantom consists of two static objects and a square moving upward from the bottom to the top at a constant speed.
@@
4.4 XCAT phantom
A simulated thoracic phantom generated by the 4DXCAT software [59] is used to assess our method on more complex structures with different motion magnitudes. This phantom has been made publicly available in [60] [^1] https://rdr.ucl.ac.uk/articles/dataset/4DCT_XCAT_phantom_dataset_for_Resolving_Variable_Respiratory_Motion_From_Unsorted_4D_Computed_Tomography_MICCAI2024/26132077. The phantom is a 3D+time volume with 182 frames and $355\times 280\times 115$ spatial voxels spanning 18 respiratory cycles. The phantom is zero padded to get $355\times 355\times 115$ voxels.
A natural question is the effect of the magnitude of the motion on the reconstruction quality. To assess this, the XCAT phantom is employed to generate 5 phantoms with different motions as follows: i) we select the first 5, 10, 20, 35, and 50 frames from the original XCAT phantom, ii) use cubic interpolation in time to get $N_{T}=100$ frames for each phantom, and iii) select the slice $z=40$ to get a 2D+time reconstruction problem. We refer to these phantoms as XCAT-$j$, for $j=5,10,20,35,50$. Each respiratory cycle lasts approximately 10 frames of the original phantom, thus XCAT-5 represents half of a cycle, while XCAT-50 represents 5 cycles. Intuitively, the reconstructed image is expected to worsen as the motion increases. An additional challenge of this phantom is that it presents out-of-plane motion with the diaphragm coming in and out of the slice, meaning that the optical flow equation is not satisfied there. Figure 6 depicts one respiratory cycle of the XCAT-5 and XCAT-50 phantoms.
We now proceed to specify the projection geometry for these phantoms. We set the distance source-origin to 6, the distance source-detector to 8, and the detector size to 3.5. We consider $N_{T}=100$ frames, and a camera with $M=150$ sensors, leading to a measurement array of dimension $150\times 100$. Measurements are further corrupted with Gaussian noise with standard deviation $0.005$. During reconstruction, each frame is generated by evaluating the neural field at a spatial grid of resolution $150\times 150$.
Figure 6: Top: XCAT-5 ground truth image at frames 10, 30, 50, 70, 90 (out of 100). XCAT-5 represents almost one respiratory cycle. Bottom: XCAT-50 ground truth image at frames 5, 10, 15, 20, 25 (out of 100). XCAT-50 represents almost five respiratory cycles.
@@ 5 Numerical Experiments
As mentioned before, in our numerical experiments, we use Tomosipo to compute the X-ray transform and its transpose. For both $u_{\theta}$ and $\mathbb{v}_{\phi}$ we use $m_{x}=m_{t}=32$ for the Fourier mappings (8). Hence, both neural fields embed $(x,t)$ into a vector of size $m=128$. We then set $L=3$ hidden layers with $d_{l}=128$ neurons each. We use $\tanh$ as activation function. Notice that we do not apply an activation function in the last layer, in particular, we do not impose the network $u_{\theta}$ to output non-negative values.
@@
We find $\sigma_{x}=\sigma_{t}=0.1$ to give the best results for the two-square and the STEMPO datasets, $\sigma_{x}=0.1,\sigma_{t}=0.5$ for the cardiac dataset, and $\sigma_{x}=\sigma_{t}=0.5$ for the XCAT-$j$ phantoms. We try two batch settings, the full-batch $N_{B}=N_{T}$ and the mini-batch $N_{B}=1$. In the full-batch setting we optimize for $150,000$ iterations, while in the mini-batch setting, we optimize for $10,000$ epochs (leading to $10,000\times N_{T}$ iterations). For the collocation points, we use a Latin Hypercube Sampling strategy [61]. We set the sampling rate (9) at $SR=0.1$, for the two-square, cardiac, and STEMPO datasets, and $SR=0.01$ for the XCAT-$j$ datasets, leading to $40,960$, $122,880$, $44,100$, and $22,500$ collocation points being randomly sampled in each iteration, respectively. We use the Adam optimizer with a learning rate of $10^{-3}$ in all cases. For the grid-based approach, each subproblem, (11) and (12), are run for $2,000$ iterations and this alternation is repeated 5 times. The PSNR values reported in this section are computed using the resolution during reconstruction. Since the ground truth is defined at a higher resolution in space, we downsample it using a pooling average. Experiments are performed on an Nvidia Tesla V100 GPU with 16GB [62]. In this paper, we are not concerned with the computational speed of the methods. Note that our naive implementation takes around a few hours for each data set (see figure 9).
5.1 Regularization of neural fields
@@
Figure 7: Ablation study on $\gamma$ for neural fields. Each plot shows the evolution of PSNR during optimization for $\gamma\in\{10^{-4},10^{-3},10^{-2},10^{-1},1\}$ and with batch size $N_{B}=N_{T}$. $\gamma=10^{-2}$ achieves superior performance in the three phantoms.
@@
Figure 8: Evolution of PSNR during optimization for the unregularized ($\gamma=0$) and the proposed regularized solution $(\gamma=10^{-2})$ with batch size $N_{B}=N_{T}$. The proposed regularized solution achieves a higher PSNR and avoids fitting noise.
@@
Figure 9: Comparing training in terms of batch size $N_{B}\in\{1,N_{T}\}$ and $\gamma\in\{0,10^{-2}\}$. Each plot shows the evolution of PSNR during optimization against time in hours. $N_{B}=1$ achieves a higher PSNR in less time but with high variability.
@@
Figure 10: Reconstruction (left) with neural fields and error (right) of the two-square phantom for the random sampling strategy at frames 35, 55, and 75 (out of 100). $\gamma=10^{-2}$ achieves smaller errors and $N_{B}=1$ gets sharper edges.
@@
In this section, we study the effect of optimizing the neural field with an explicit motion regularizer and compare it against the purely implicitly regularized solution. For this purpose, we consider a simplification of the variational problem (5) by setting the regularization parameters $\alpha,\beta=0$ and performing an ablation study on $\gamma\in\{10^{-4},10^{-3},10^{-2},10^{-1},1\}$ in the two-square, cardiac, and STEMPO datasets with random sampling. This choice helps to understand the role of the optical flow by isolating it from the other regularizers $\mathcal{R}$ and $\mathcal{S}$.
@@
We also compare the full-batch setting $N_{B}=N_{T}$ against the commonly used mini-batch setting with $N_{B}=1$, and for $\gamma\in\{0,10^{-2}\}$. We recall that for $N_{B}=1$ optimization is done for 10,000 epochs. Results for the evolution of PSNR versus time are shown in figure 9. A batch size of 1 speeds up the optimization, allowing to achieve a higher PSNR in less time. It suffers however from an unstable behaviour with the PSNR varying for almost 2 points every 100 epochs. This means we can stop at a poor reconstruction without a reliable stopping criteria.
Obtained reconstructions at their best PSNR and the corresponding error against the ground truth at different frames for the two-square, cardiac, and STEMPO datasets are displayed in figures 10, 12, and 12, respectively. The top rows show the full-batch regularized case $N_{B}=N_{T},\gamma=10^{-2}$; the middle rows show the mini-batch regularized case $N_{B}=1,\gamma=10^{-2}$; the bottom rows show the full-batch unregularized case $N_{B}=N_{T},\gamma=0$. A relevant advantage in the mini-batch setting is that it is more likely to capture edges in less time since more iterations can be taken, while its full-batch counterpart $N_{B}=N_{T}$ shows blurry edges. We also summarize the PSNR values obtained for the different models in table 1, highlighting the role of the PDE-based regularizer.
@@
Figure 11: Reconstruction (left) with neural fields and error (right) of the cardiac phantom for the random sampling strategy at frames 1, 122, and 250 (out of 300). $\gamma=10^{-2}$ achieves smaller errors and $N_{B}=1$ gets sharper edges.
Figure 12: Reconstruction (left) with neural fields and error (right) of the STEMPO phantom for the sequential 32-degree sampling strategy at frames 1, 40, and 60 (out of 90). $\gamma=0$ gets larger errors.
We finish this section by comparing the PDE-based motion regularizer against the more common total variation and show the benefits of employing the former. To showcase the effects of the TV regularizer, we proceed in a similar way as in the previous experiment, this is, setting $\beta,\gamma=0$ and performing an ablation study on $\alpha$. Additionally, to complement our study, we consider a third regularizer, the spatiotemporal total variation (STV) regularizer, which differs from $\mathcal{R}$ in that it penalizes changes in time as well and is defined as follows:
^$\hat{\mathcal{R}}(u):=\displaystyle\int_{\Omega_{T}}\sqrt{\|\nabla u\|^{2}+(\partial_{t}u)^{2}}.$
We consider the two-square phantom for this experiment. Results displayed in figure 13 show that the best reconstruction using the optical flow constraint enhances the reconstruction task by almost 5 when compared to the best reconstruction using TV or STV regularization.
Figure 13: Comparing optical flow and TV-like regularizers on the two-square phantom. Left: ablation study on $\gamma$ while setting $\alpha,\beta=0$. Center: ablation study on $\alpha$ while setting $\beta,\gamma=0$ using the TV regularizer in space. Right: ablation study on $\alpha$ while setting $\beta,\gamma=0$ using the STV regularizer. A higher PSNR is attained for the optical flow regularizer.
5.2 Effects of rapid motion
In this section we employ the XCAT-$j$ phantoms to study the effects of motion in the reconstruction. We recall that, by construction, the larger is $j$, the larger is the motion. For this experiment we set the Fourier feature hyperparameters $\sigma_{x}=\sigma_{t}=0.5$, the regularization parameters $(\alpha,\beta,\gamma)=(10^{-5},10^{-5},10^{-3})$, the batch size $N_{B}=1$, and train for 10,000 epochs for the five XCAT-$j$ phantoms.
Figure 14 shows reconstructions and the corresponding errors against the ground truth image for XCAT-5 and XCAT-50 phantoms at comparable frames that span part of one respiratory cycle. As expected, a larger error is obtained for XCAT-50. This is supported in figure 15 which shows the PSNR achieved for each phantom after optimization finished and reveals an almost linear decay on the reconstruction with respect to the velocity. We also notice a difficulty in capturing fine details given by the tiny dots (pulmonary alveoli) within the lungs.
The circular structure of the scene with varying sizes is the diaphragm. This structure represents out-of-plane motion, which clearly violates the brightness constancy assumption imposed by the optical flow model. However, given that this is imposed as a soft constraint in the variational problem, our method can still get a reliable reconstruction through the data-fidelity term.
Figure 14: Reconstruction (left) with neural fields and error (right) for XCAT-5 at frames 50, 70, 90 and XCAT-50 at frames 5, 7, 8. Depicted frames for XCAT-5 and XCAT-50 represent similar time instances from the first respiratory cycle. Motion magnitude is larger in XCAT-50, thus, larger errors are obtained.
Figure 15: PSNR between reconstruction and ground truth for XCAT-$j$ phantoms, with $j=5,10,20,35,50$. The motion in XCAT-$j$ increases with $j$, hence, reconstruction quality decreases.
5.3 Regularized neural fields versus grid-based method
5.3.1 Random sampling
Figure 16: Neural field versus grid-based for the two-square phantom. Neural fields achieve smaller errors when compared to the grid-based solution.
Figure 17: Neural field versus grid-based for the cardiac phantom. The grid-based method fails at the reconstruction of the cardiac phantom while neural fields can capture the intricate motion.
@@ 5.3.1 Random sampling
Figure 18: Comparing grid-based method against neural fields for the two-square phantom with sequential sampling. Reconstruction (left) and error (right) using the regularization parameters that led to the best results for random sampling. Neural fields still perform better than grid-based but both reconstructions worsen with respect to the one obtained with random sampling.
Figure 19: Comparing grid-based method against neural fields for the STEMPO phantom with sequential 4-degree sampling. Reconstruction (left) and error (right). Both the neural field and the grid-based solutions fail at the reconstruction task. Neural field Adaptive increases the value of $\gamma$ to improve the reconstruction.
We now compare explicitly regularized neural fields against the grid-based method outlined in section 3.2 for the random sampling regime. For this approach, it was found that the choice $(\alpha,\beta,\gamma)=(10^{-3},10^{-4},10^{-3})$ led to the best results in terms of PSNR, achieving a value of 27.09. We solve the same variational problem with neural fields by choosing the same regularization parameters, in which case the PSNR achieved is 32.92. Figure 16(a) shows reconstructions at different frames. To appreciate the time evolution we also show a $x-t$ slice view in figure 16(b) at the horizontal center $y=0$, where it becomes clear that the grid-based solution struggles at capturing regularity in time.
@@
5.3.2 Sequential sampling
We finish our study by comparing both methods at the more challenging problem of sequential sampling. We try the sequential 9-degree sampling for the two-square phantom and the 4-degree sampling for STEMPO. For the two-square phantom, we use the regularization parameters that gave the best results in the previous section, namely, $(\alpha,\beta,\gamma)=(10^{-3},10^{-4},10^{-3})$ for the grid-based method achieving a PSNR of 22.65, while for the neural field, we use $(\alpha,\beta,\gamma)=(0,0,10^{-2})$ and the method achieves a PSNR of 26.42. Results are shown in figure 19. There it can be seen that both methods perform worse than in the random sampling, but still, the neural field reconstruction is better than the grid-based one which shows large errors, for instance, in frame 60. For STEMPO we found $(\alpha,\beta,\gamma)=(10^{-4},10^{-4},10^{-1})$ to give the best reconstruction for the grid-based method with a PSNR of 14.24; for the neural field, we set $(\alpha,\beta,\gamma)=(0,0,10^{-2})$, leading to a PSNR of 15.18. Results can be seen in figure 19. In this case, both methods can capture the static part of the image but fail at the reconstruction of the moving square due to the sampling scheme.
@@ 5.3.2 Sequential sampling
6 Discussion
Despite neural fields being regarded as resolution-independent representations, they still need to be queried at spatiotemporal grid coordinates to evaluate the forward operator. Since the evaluation of each coordinate requires a forward pass of the network, this process may become computationally inefficient and time-consuming for large-scale problems (see, for instance, the runtimes in figure 9). Thus, a limitation of the spatiotemporal neural field used in this work arises when scaling to a 3D+time scenario with demand for high spatial resolution, as in dynamic cone-beam CT.
This limitation can be addressed by novel positional encodings, such as hash enconding [63] or ACORN [37]. These encodings work at multiscale resolution levels and can capture fine details at a lower computational cost. In particular, they resulted in faster and higher-quality reconstructions compared to the Fourier encoding used in this work. Another line of research accelerates the optimization by applying the forward operator on auxiliary image variables instead of acting on the rasterized image obtained from the neural field at every iteration [41, 64]. In such approaches, weight updates occur entirely in the image domain. Importantly, the PDE-based regularizer employed in this work can be seamlessly integrated into such frameworks, highlighting its potential when scaling to more realistic and computationally demanding settings.
@@ 7 Conclusion
This work considers neural fields for dynamic CT image reconstruction from finely resolved in time but severely undersampled in space measurements as commonly encountered in applications, e.g. cardiac CT. Neural fields are particularly suitable for this task: their continuity allows for a coherent representation in both time and space; their resolution independence enables us to query it at multiple frames, and, together with their differentiability, facilitate the numerical evaluation of the PDE-based motion regularizer. We studied how to enhance the neural field reconstruction by making use of the optical flow equation: constraining the neural field to this physically feasible motion meant a significant improvement with respect to the state-of-the-art both in terms of neural fields as well as traditional grid-based representations. We finish by mentioning that further regularization techniques for dynamic imaging with neural fields can be studied, such as sparsity in suitable domains or deep-learning-based ones. This is left for future research.